OpenStudy (adamaero):

MEDAL How to make an integral with these bounds? pi/6 <= y <= y^2

OpenStudy (adamaero):

\[\frac{ \pi }{ 6 }\le y \le y^2\]

OpenStudy (adamaero):

@ganeshie8

OpenStudy (delirium):

tricks-y, lol

OpenStudy (delirium):

I could probably get you the answer using a calculator, but I'm not sure HOW to get there myself...